# How to Build a Quantum Supercomputer: Scaling Challenges and Opportunities

### v1

@@ Abstract
In the span of four decades, quantum computation has evolved from an intellectual curiosity to a potentially realizable technology. Today, small-scale demonstrations have become possible for quantum algorithmic primitives on hundreds of physical qubits and proof-of-principle error-correction procedures on a single logical qubit. Nevertheless, despite significant progress and excitement, the detailed path toward a full-stack scalable quantum computing technology is largely unknown. There are significant outstanding quantum hardware, fabrication, software architecture, and algorithmic challenges that are either unresolved or overlooked. These issues could seriously undermine the arrival of utility-scale quantum computers for the foreseeable future. Here, we provide a comprehensive review of these scaling challenges. We show how the road to scaling could be paved by adopting existing semiconductor technology to build much higher-quality qubits, employing system engineering approaches, and performing distributed quantum computation within heterogeneous high-performance computing infrastructures. These opportunities for research and development could unlock certain promising applications, in particular, efficient quantum simulation and learning/modeling quantum data generated by natural or engineered quantum systems. In order to estimate the true cost of such promises, we provide a detailed resource and sensitivity analysis for classically hard quantum chemistry calculations on surface-code error-corrected quantum computers given current, target, and desired hardware specifications based on superconducting qubits, accounting for a realistic distribution of errors. We show orders of magnitude enhancement in performance could be obtained by a combination of quantum hardware and algorithmic improvements. Furthermore, we argue that, to tackle today’s industry-scale classical optimization and machine learning problems in a cost-effective manner, distributed quantum-assisted probabilistic computing with custom-designed accelerators should be considered as a complementary path toward scalability.
@@ I Introduction: Past, present, and future challenges of building quantum computers
Quantum computing has seen remarkable progress over the past few decades despite facing tremendous conceptual, theoretical, and technical challenges. There have been a few significant breakthroughs, such as Shor’s algorithm for integer factoring Shor (1994) to solve a seemingly exponentially hard classical problem, or the invention of quantum error-correction (QEC) Shor (1995a); Gottesman (1997) to tackle the fundamental problem of decoherence, without which Shor’s algorithm would remain merely a mathematical curiosity. Starting from small experiments that manipulate single- or few-qubit systems, research groups using a variety of technologies can now make and operate quantum processors with on the order of 100 physical qubits. Some proof-of-principle speedups over conventional (classical) supercomputers called “quantum supremacy” Arute et al. (2019) or “quantum advantage” qua (2023) have been demonstrated, but only for carefully crafted problems. The next step is to scale up quantum processors to demonstrate significant speedup for a practical problem in a cost effective manner, thereby achieving “quantum utility” Altepeter (2022).
@@
Recently, there has been a significant interest in the high-performance computing (HPC) community in employing quantum computation as a complementary paradigm beyond exascale supercomputing Alexeev (2024); Bravyi et al. (2022); Humble et al. (2021). Rather than replacing classical computers as general-purpose processors, quantum computers can be better understood as accelerators or coprocessors that can efficiently carry out specialized tasks within an HPC framework. Hybrid quantum–classical frameworks will be crucial not only in the near term—the noisy, intermediate scale quantum (NISQ) era Preskill (2018)—but also for future fault-tolerant quantum computation (FTQC), as error-correction schemes will rely heavily on classical HPC and the number of logical qubits will be fairly small for the foreseeable future. To achieve true utility-scale quantum computing, successful integration with existing heterogeneous HPC infrastructures and the development of a hybrid quantum–classical full computing stack are necessary.
@@
There are a number of challenges to successful quantum-HPC integration. At the hardware architecture and system design levels, there are significant differences between the quantum and classical components in physical scale, hardware reliability, control electronics, communication bandwidth, and time scales of operation. At the algorithmic level, the challenges lie within memory access, data sharing and movement, and information extraction. To build a quantum-centric supercomputer Alexeev (2024); Bravyi et al. (2022) at scale, much better quantum components must be built at different layers that ultimately rely on much higher quality qubits. While basic research remains critical, a comprehensive engineering approach targeting the full stack must be taken in parallel, with the aim of steadily increasing the technology readiness level (TRL) of the components at various levels of the full stack. This mindset is needed for NISQ devices, logical intermediate-scale quantum (LISQ) processors, and for FTQC.
@@
Utility-scale quantum computing ultimately involves very deep quantum circuits requiring logical error rates far below those experimentally feasible with physical qubits and gate operations. Quantum error correction Shor (1995a); Gottesman (1997) is necessary, as even very small errors rapidly accumulate, resulting in significant errors. Similar to classical error correction, QEC utilizes redundancy in the encoding of information to detect and correct errors. However, unlike in the case of classical telecommunication, copying quantum information is prohibited due to the no-cloning theorem Wootters and Zurek (1982). In addition, quantum measurements irreversibly collapse the wavefunction. Therefore, the key mechanism for error correction without destroying logical information is through exploiting quantum entanglement with ancillary subsystems. Quantum error-correction codes (QECC) protect a smaller logical state of computation within a much larger highly entangled quantum state involving many noisy physical qubits. Therefore, errors below a certain weight can be detected and corrected. Multi-qubit stabilizer measurements produce a set of syndromes that detect whether some of the physical qubits have been corrupted. A QEC decoder can then infer the most likely errors for the detected syndromes. With this knowledge, corrective operations can be either physically applied (active error correction) or kept track of in software.
@@
While QEC is necessary, it is by no means sufficient for reliable large-scale quantum computation. When performing QEC, the operations on the physical qubits necessary to implement the QECC can eventually themselves add more errors than they are able to correct. Thus, achieving FTQC requires implementation schemes in which QEC succeeds in suppressing errors faster than it causes them. According to the threshold theorem, once the physical errors of the quantum system are below a certain threshold, the overhead of these schemes scales poly-logarithmically with the precision of computation Aharonov and Ben-Or (1997); Knill et al. (1998); Kitaev (2003a). However, these schemes introduce a substantial overhead in physical resources, the estimating and validating of which is a critical step toward practical realization of FTQC at utility scale and the cost of doing so.
Here, at the lowest level of the full stack—the physical components for storing and processing quantum information—we focus on superconducting qubits. Historically, the success of superconducting qubits has come from a sustained focus on improving qubit coherence using different approaches. At present, although there are a variety of ideas on how to make further advances, there is sufficient justification for using advanced semiconductor processing and tools. This approach will need to be guided by and integrated with more complex device architectures and structures, such as bump-bonding and advanced packaging, which are designed around specific decoherence mechanisms such as large two-level state loss from amorphous insulators.
@@
These advances were also shaped by an understanding that qubit architecture would be improved by designs that turned off the qubit–qubit interaction, which does not occur naturally using simple coupling capacitors. The adjustable coupler was pioneered by UCSB and Google and is an example of careful systems engineering that was initially assessed skeptically by a majority of the superconducting qubit community. Although the adjustable coupler was significantly more complex, as it required using a qubit to turn interactions on and off, the reduction in crosstalk enabled Google to achieve the quantum supremacy milestone Arute et al. (2019). Today, the adjustable coupler has been adopted and integrated into superconducting systems from IBM, USTC, Rigetti, and IQM. Alternative superconducting qubit architectures such as the fluxonium are only being realized Ding et al. (2023) thanks to the adjustable couplers technology. Further innovations which, like adjustable couplers, trade simplicity for performance are needed to push errors in large qubit systems to the $10^{-4}$ range.
@@
Despite the large size of superconducting qubits, roughly millimeters, the numbers of qubits have been scaled up in a brute-force manner once single- and few-qubit systems were established. Indeed, recent advances enabled building large enough quantum computers to test system performance and demonstrate the execution of algorithms. Experiments have even achieved logical error rates at the $10^{-10}$ level, although surpassing this rate appears to be obstructed by errors induced via cosmic rays McEwen et al. (2024), which demands further investigation. Although it has been speculated that the large footprint of superconducting qubits makes them too difficult to scale, in this position paper we describe how modern semiconductor processing may be used to solve this problem.
@@
While we introduce a definite architecture for coupling transmon qubits, alternative designs might also improve performance. Our expectation is that advanced fabrication targeted to improve both coherence and scaling could be used for a variety of approaches. We believe these ideas will greatly enhance the likelihood of making useful quantum computers.
@@ I.1 Systems engineering for quantum hardware
To make these advances, the most important concept is systems engineering, where one embraces the idea that many system parameters must be simultaneously optimized for a complex system. A significant obstacle in prior research stems from treating quantum computation as a highly tailored quantum physics experiment, with physicists naturally focusing on isolating physical phenomena in order to fully understand each of them. For example, tables of metrics often emphasize the best quality of the various approaches, whereas system engineering typically constrains the system with its worst quality.
There are many systems engineering parameters that must be considered and they differ for each technology. Here, we focus on the four most important parameters that serve as a concise but powerful way to compare quantum hardware approaches.
@@
Quantity. This is the most intuitively understood metric. As discussed above, at first we need to optimize for low error rates. However, as errors reach the 0.01–0.1% range, the ability to scale up the number of qubits becomes more important. This is, in part, because qubit number scales logarithmically with the error suppression rate of QECCs such as the surface code. So, once the error rate is sufficiently low, the strategy should change to prioritizing qubit counts. Scaling to thousands or millions of qubits is required in the long term, and thus careful deliberation is needed on how to achieve that with any particular technology.
@@
Connectivity. The number of connections from one physical qubit to another is also an important metric, and has an interesting trade-off with speed. For neutral atom and trapped-ion systems, the connectivity is generally considered all-to-all, but this may be the case only for small enough systems, e.g., within a single trap. This all-to-all connectivity also typically requires ample time for shuttling qubits. Superconducting qubits do not need to move but have sparse connectivity; they typically have nearest-neighbor interactions, either to 4 qubits in a regular rectangular lattice or 2.5 qubits on average for the heavy-hex lattice ibm (2021). This more-limited connectivity can be factored into the design of near-term algorithms, and thus it is hard to make a fair performance comparison with respect to connectivity. For a fault-tolerant quantum computer, the present connectivity of superconducting qubits is sufficient to support error-corrected logical qubits such as the surface codes. Additionally, the more-connected rectangular lattice architecture is perceived to be less degraded by qubit dropouts, an important system constraint.
@@ I.2 Technical challenges and opportunities at different scales
Ultimately, efforts to scale the number of physical and logical qubits in quantum processors must rely on QECCs. We foresee that different scales, and possibly different applications, may favor different types and sizes of error-correcting codes. Even in the case of surface codes, different variants (e.g., the XZZX or XY codes Bonilla Ataides et al. (2021); Tsai et al. (2024)) may be preferable for different hardware noise profiles. These codes must be supported by architectural provisions for fast classical decoding and control based on syndrome measurement results. Additional support is required at the operating and compilation level.
@@ I.2.1 Challenges at 100–1000 physical qubits
At the intermediate scale, key challenges involve individual qubits and gates operating within the system, as well as fabrication and basic operation Markov (2014). At this scale, executing quantum algorithms is used primarily to demonstrate and characterize hardware capabilities.
@@
Circuit knitting overhead. In the past few years, circuit knitting methods have been introduced to allow running quantum circuits that require more qubits than are available on a single processor at the current scale Peng et al. (2020); Piveteau et al. (2022); Piveteau and Sutter (2023). Formally, these methods incur an exponential classical post processing overhead for exact reconstruction of a quantum observable. To enable distributed quantum circuit execution at scale, innovative techniques must be devised for reducing this exponential overhead. While the challenge of quantum workload distribution emerges at the scale of 100-1000 physical qubits, it will be present at all later scales. In Section V.2 we provide a detailed discussion on this topic as well as present a family of adaptive circuit knitting methods. In Section VI.2, we show a particular implementation of this adaptive circuit knitting approach that could significantly reduce overheads for quantum simulation of quantum spin glasses via an approximate tensor-network contraction over distributed quantum circuits.
@@
We also note that, in the near-term with qubit counts below 1000, automated testing techniques at the system and component level are necessary. We discuss these procedures in Section II.1 on qubit fabrication for component-level testing.
@@ I.2.2 Challenges at 1000–10k physical qubits
The largest risk of this phase is the cost of development of these processes, which could be mitigated by leveraging the semiconductor industry. In Section II.2 on wafer-scale integration, we also discuss how to leverage the existing semiconductor industry to drastically reduce costs (see also Section VIII.1). Because of the high costs of developing, building, and operating fault-tolerant quantum computers, there must be a strong emphasis on understanding the impact of exact hardware noise profiles on the choice of error correcting codes, as well as the resulting resource estimates for useful applications with utility-scale value. In Section III.2 we explain how we use hardware noise profiles at this scale to inform FTQC compilation and assembly at the utility scale.

### v2

# How to Build a Quantum Supercomputer: Scaling from Hundreds to Millions of Qubits

@@ Abstract
In the span of four decades, quantum computation has evolved from an intellectual curiosity to a potentially realizable technology. Today, small-scale demonstrations have become possible for quantum algorithmic primitives on hundreds of physical qubits and proof-of-principle error-correction procedures on a single logical qubit. Nevertheless, despite significant progress and excitement, the detailed path toward a full-stack scalable quantum computing technology is largely unknown. There are significant outstanding quantum hardware, fabrication, software architecture, and algorithmic challenges that are either unresolved or overlooked. These issues could seriously undermine the arrival of utility-scale quantum computers for the foreseeable future. Here, we provide a comprehensive review of these scaling challenges. We show how the road to scaling could be paved by adopting existing semiconductor technology to build much higher-quality qubits, employing systems engineering approaches, and performing distributed quantum computation within heterogeneous high-performance computing infrastructures. These opportunities for research and development could unlock certain promising applications, in particular, efficient quantum simulation and learning quantum data generated by natural or engineered quantum systems. In order to estimate the true cost of such promises, we provide a detailed resource and sensitivity analysis for classically hard quantum chemistry calculations on surface-code error-corrected quantum computers given current, target, and desired hardware specifications based on superconducting qubits, accounting for a realistic distribution of errors. We show that orders of magnitude enhancement in performance could be obtained by a combination of quantum hardware and algorithmic improvements. Furthermore, we argue that, to tackle today’s industry-scale classical optimization and machine learning problems in a cost-effective manner, heterogeneous quantum-probabilistic computing with custom-designed accelerators should be considered as a complementary path toward scalability.
@@ I Introduction: Past, present, and future challenges of building quantum computers
Quantum computing has seen remarkable progress over the past few decades despite facing tremendous conceptual, theoretical, and technical challenges. There have been a few significant breakthroughs, such as Shor’s algorithm for integer factoring Shor (1994) to solve a seemingly exponentially hard classical problem, or the invention of quantum error-correction (QEC) Shor (1995a); Gottesman (1997) to tackle the fundamental problem of decoherence, without which Shor’s algorithm would remain merely a mathematical curiosity. Starting from small experiments that manipulate single- or few-qubit systems, research groups using a variety of technologies can now make and operate quantum processors with on the order of 100 physical qubits. Some proof-of-principle speedups over conventional (classical) supercomputers called “quantum supremacy” Arute et al. (2019) or “quantum advantage” qua (2023) have been demonstrated, but only for carefully crafted problems. The next step is to scale up quantum processors to demonstrate significant speedup for a practical problem in a cost effective manner, thereby achieving quantum utility Altepeter (2022).
@@
Recently, the high-performance computing (HPC) community has shown significant interest in employing quantum computation as a complementary paradigm beyond exascale supercomputing Alexeev (2024); Bravyi et al. (2022); Humble et al. (2021). Rather than replacing classical computers as general-purpose processors, quantum computers can be better understood as accelerators or coprocessors that can efficiently carry out specialized tasks within an HPC framework. Hybrid quantum–classical frameworks will be crucial not only in the near term—the noisy, intermediate-scale quantum (NISQ) era Preskill (2018)—but also for future fault-tolerant quantum computation (FTQC), as error-correction schemes will rely heavily on classical HPC and the number of logical qubits will be fairly small for the foreseeable future. To achieve true utility-scale quantum computing, successful integration with existing heterogeneous HPC infrastructures and the development of a hybrid quantum–classical full computing stack are necessary.
@@
Practical quantum–HPC integration runs into a number of challenges. At the hardware-architecture and system-design levels, there are significant differences between the quantum and classical components in physical scale, hardware reliability, control electronics, communication bandwidth, and time scales of operation. At the algorithmic level, the challenges lie within memory access, data sharing and movement, and information extraction. To build a quantum-centric supercomputer Alexeev (2024); Bravyi et al. (2022) at scale, much better quantum components must be built at different layers that ultimately rely on much higher quality qubits. While basic research remains critical, a comprehensive engineering approach targeting the full stack must be taken in parallel, with the aim of steadily increasing the technology readiness level (TRL) of the components at various levels of the full stack. This mindset is needed for NISQ devices, logical intermediate-scale quantum (LISQ) processors, and for FTQC.
@@
Utility-scale quantum computing ultimately involves deep quantum circuits requiring logical error rates far below those experimentally feasible with physical qubits and gate operations. Quantum error correction Shor (1995a); Gottesman (1997) is necessary, as even small errors rapidly accumulate, resulting in significant errors. Similar to classical error correction, QEC utilizes redundancy in the encoding of information to detect and correct errors. However, unlike in the case of classical telecommunication, copying quantum information is prohibited due to the no-cloning theorem Wootters and Zurek (1982). In addition, quantum measurements irreversibly collapse the wavefunction. Therefore, the key mechanism for error correction without destroying logical information is through exploiting quantum entanglement with ancillary subsystems. Quantum error-correction codes (QECC) protect a smaller logical state of computation within a much larger highly entangled quantum state involving many noisy physical qubits. Therefore, errors below a certain weight can be detected and corrected. Multi-qubit stabilizer measurements produce a set of syndromes that detect whether some of the physical qubits have been corrupted. A QEC decoder can then infer the most likely errors for the detected syndromes. With this knowledge, corrective operations can be either physically applied (active error correction) or kept track of in software.
@@
While QEC is necessary, it is by no means sufficient for reliable large-scale quantum computation. When performing QEC, the operations on the physical qubits necessary to implement the QECC can eventually themselves add more errors than they are able to correct. Thus, achieving FTQC requires implementation schemes in which QEC succeeds in suppressing errors faster than it causes them. According to the threshold theorem, once the physical errors of the quantum system are below a certain threshold, the overhead of these schemes scales poly-logarithmically with the precision of computation Aharonov and Ben-Or (1997); Knill et al. (1998); Kitaev (2003a). However, these schemes introduce substantial overhead in physical resources, whose estimation and validation is a critical step toward practical, cost-efficient realization of FTQC at utility scale.
@@
At the lowest level of the full stack — the physical components for storing and processing quantum information — we focus on superconducting qubits. Historically, the success of superconducting qubits has come from a sustained focus on improving qubit coherence using different approaches. At present, although there are a variety of ideas on how to make further advances, there is sufficient justification for using advanced semiconductor processing and tools. This approach will need to be guided by and integrated with more complex device architectures and structures, such as bump-bonding and advanced packaging, which are designed around specific decoherence mechanisms such as large two-level state loss from amorphous insulators. These advances were also shaped by an understanding that qubit architecture would be improved by designs that turned off the qubit–qubit interaction, which does not occur naturally using simple coupling capacitors. The adjustable coupler was pioneered by UCSB and Google and is an example of careful systems engineering that was initially assessed skeptically by a majority of the superconducting qubit community. Although the adjustable coupler was significantly more complex, as it required using a qubit to turn interactions on and off, the reduction in crosstalk enabled Google to achieve the quantum supremacy milestone Arute et al. (2019). Today, the adjustable coupler has been adopted and integrated into superconducting systems from IBM, USTC, Rigetti, and IQM. Alternative superconducting qubit architectures such as the fluxonium are only being realized Ding et al. (2023) thanks to the adjustable couplers technology. Further innovations which, like adjustable couplers, trade simplicity for performance are needed to push errors in large qubit systems to the $10^{-4}$ range.
@@
Despite the large size of superconducting qubits (roughly millimeters), the numbers of qubits have been scaled up from single- and few-qubit systems. Recent advances enabled large enough quantum computers to test system performance and demonstrate the execution of simple quantum algorithms. Experiments have demonstrated logical error rates at the $10^{-10}$ level, although surpassing this rate appears difficult due to errors induced via cosmic rays McEwen et al. (2024), suggesting further investigation. Despite the large footprint of superconducting qubits, in this position paper we describe how modern semiconductor processing facilitates their scaling.
@@
While we introduce a definitive architecture for coupling transmon qubits, alternative designs might also improve performance. Our expectation is that advanced fabrication targeted to improve both coherence and scaling could be used for a variety of approaches. We believe that these ideas will greatly enhance the likelihood of making useful quantum computers.
@@ I.1 Systems engineering for quantum hardware
Systems engineering simultaneously optimizes many system parameters. In contrast, prior research often treats quantum computation as a custom-configured quantum physics experiment focusing on isolated physical phenomena to fully understand each of them. For example, reported metrics often emphasize the best-case quality of a given approach, whereas systems engineering is typically sensitive to average and worst-case quality.
Each technology is sensitive to a different set of systems engineering parameters. Here, we focus on the four most important parameters that serve as a concise but powerful way to compare quantum hardware approaches.
@@
Quantity. This is the most intuitively understood metric. As discussed above, at first we need to optimize for low error rates. However, as errors reach the 0.01–0.1% range, the ability to scale up the number of qubits becomes more important. This is, in part, because the number of qubits scales logarithmically with the error-suppression rate of QECCs such as the surface code. So, once the error rate is sufficiently low, the strategy should change to prioritizing qubit counts. Scaling to thousands or millions of qubits is required in the long term, and thus careful deliberation is needed on how to achieve that with any particular technology.
@@
Connectivity. The number of connections from one physical qubit to another is also an important metric, and has an interesting trade-off with speed. Neutral-atom and trapped-ion systems often advertise all-to-all connectivity, but this might not scale beyond small systems, e.g., single traps. This all-to-all connectivity also typically requires ample time for shuttling qubits. Superconducting qubits do not need to move but have sparse connectivity; they typically have nearest-neighbor interactions, either to 4 qubits in a regular rectangular lattice or 2.5 qubits on average for the heavy-hex lattice ibm (2021). This more-limited connectivity can be factored into the design of near-term algorithms, and thus it is hard to make a fair performance comparison with respect to connectivity. For a fault-tolerant quantum computer, the present connectivity of superconducting qubits is sufficient to support error-corrected logical qubits such as the surface codes. Additionally, the more-connected rectangular lattice architecture is less likely to suffer from qubit dropouts, an important system constraint.
@@ I.2 Technical challenges and opportunities at different scales
Ultimately, efforts to scale the number of physical and logical qubits in quantum processors must rely on QECCs. We foresee that different scales and different applications may favor different types and sizes of error-correcting codes. Even in the case of surface codes, different variants (e.g., the XZZX or XY codes Bonilla Ataides et al. (2021); Tsai et al. (2024)) may be preferable for different hardware noise profiles. These codes must be supported by architectural provisions for fast classical decoding and control based on syndrome measurement results. Additional support is required at the operating and compilation level.
@@ I.2.1 Challenges at 100–1000 physical qubits
At the intermediate scale, key challenges involve individual qubits and gates operating within the system, as well as fabrication and basic operation Markov (2014). At this scale, the execution of quantum algorithms is used primarily to demonstrate and characterize hardware capabilities.
@@
Circuit knitting overhead. In the past few years, circuit knitting methods have been introduced to allow running quantum circuits that require more qubits than are available on a single processor at the current scale Peng et al. (2020); Piveteau et al. (2022); Piveteau and Sutter (2023). Formally, these methods incur an exponential classical post processing overhead for exact reconstruction of a quantum observable. To enable distributed quantum circuit execution at scale, innovative techniques must be devised for reducing this exponential overhead. While the challenge of quantum workload distribution emerges at the scale of 100–1000 physical qubits, it will be present at all later scales. In Section V.3 we provide a detailed discussion on this topic as well as present a family of adaptive circuit knitting methods. In Section VI.2, we show a particular implementation of this adaptive circuit knitting approach that could significantly reduce overheads for quantum simulation of quantum spin glasses via an approximate tensor-network contraction over distributed quantum circuits.
@@
We also note that, with qubit counts below 1000 in the near future , automated testing techniques at the system and component level are necessary. We discuss these procedures in Section II.1 on qubit fabrication for component-level testing.
@@ I.2.2 Challenges at 1000–10k physical qubits
The largest risk of this phase is the cost of development of these processes, which could be mitigated by leveraging the semiconductor industry. In Section II.2 on wafer-scale integration, we also discuss how to leverage the existing semiconductor industry to drastically reduce costs (see also Section VIII.2). Because of the high costs of developing, building, and operating fault-tolerant quantum computers, there must be a strong emphasis on understanding the impact of exact hardware noise profiles on the choice of error correcting codes, as well as the resulting resource estimates for useful applications with utility-scale value. In Section III.2 we explain how we use hardware noise profiles at this scale to inform FTQC compilation and assembly at the utility scale.
Cryogenic thermal and space budgets. A key systems engineering challenge is to design the components in a way that it fits within the thermal and space budgets afforded by the cryogenic system. In general, the lower the temperature stage within the cryostat is, the less cooling power and space there is available for the various components. The proposed approach carefully systems engineers the various components accounting for this constraint, balancing qubit quality, complexity of the components, and complexity of fabrication. These constraints become especially acute when scaling from 1000 to 10k qubits. Detailed power and space budgets are not provided in this paper due to proprietary designs.